# Tactile SoftHand-A: 3D-Printed, Tactile, Highly-underactuated, Anthropomorphic Robot Hand with an Antagonistic Tendon Mechanism

### v1

@@ 10 Conclusion
The experimental results provide evidence of the advanced control mechanisms and adaptability of the Tactile SoftHand-A. The coordinated control of driving and antagonistic motors allows for precise manipulation of finger joints, while the differential mechanism enhances grasping adaptability. The tactile feedback system further improves the hand’s responsiveness and stability in real-world scenarios. These findings contribute to the development of more dexterous and functional minimally-actuated tendon-driven robot hands, capable of performing a wide range of tasks with precision and reliability. Future work could focus on refining these control mechanisms and exploring their applications in dexterous manipulation tasks. Further, because the hand is 3D-printed, different combinations of driving and antagonistic mechanisms coupled with the degree of actuation can be explored. This opens up a rich research area of adding greater dexterity to minimally-actuated anthropomorphic robot hands, so that it becomes possible to investigate systematically how to develop low-cost, minimally actuated hands with evermore human-like dexterity. To this end, we openly release all designs and fabrication instructions for creating the Tactile SoftHand-A as an enabler for others to build upon this work.
@@ 11 Funding
HL, CL and YL were supported by China Scholarship Council (CSC) and Bristol joint scholarship. NFL was supported by a research leadership award from the Leverhulme Trust: ‘A Biomimetic Forebrain for Robot Touch’ (RL-2016-039). CJF, EP and NFL were supported by the European Union’s Horizon Europe program under Grant Agreement No 101120823 (’MANiBOT: Advancing the physical intelligence and performance of roBOTs towards human-like bi-manual objects MANipulation’).

### v2

@@ 10 Conclusion
This paper shows the sophisticated control and adaptivity possible with a 3D-printed anthropomorphic SoftHand with an antagonist tendon mechanism and integrated tactile sensing. The coordinated control of agonist and antagonist motors allows for precise manipulation of finger joints, while the differential mechanism enhances grasping adaptivity. The tactile feedback system further improves the hand’s responsiveness and stability in autonomous and teleoperation scenarios. These findings have implications for the application of tendon-driven dexterous hands, for example as prosthetics for assisting individuals with disabilities to perform daily grasping tasks and for remote operations in hazardous environments or precision surgeries to enable more accurate and reliable human-machine interactions. Future work could focus on refining these control mechanisms and exploring their applications in dexterous manipulation. Further, because the hand is 3D-printed, different combinations of agonist and antagonistic mechanisms coupled with the degree of actuation can be explored. This opens up a rich research area of adding greater dexterity to minimally-actuated anthropomorphic robot hands, so that it becomes possible to investigate systematically how to develop low-cost hands with human-like dexterity. To this end, we openly release all designs and fabrication instructions for creating the Tactile SoftHand-A as an enabler for others to build upon this work.
@@ 11 Funding
CL, HL and YL were supported by the China Scholarship Council (CSC) and Bristol joint scholarship. EP and NL were supported by the Horizon Europe research and innovation program under grant agreement No. 101120823 (‘MANiBOT: Advancing the physical intelligence and performance of roBOTs towards human-like bi-manual objects MANipulation’) and a Royal Society International Collaboration Award (South Korea). NL was also supported by the Advanced Research + Invention Agency (‘Democratising Hardware and Control For Robot Dexterity’).